# DebiasDiff: Debiasing Text-to-image Diffusion Models with Self-discovering Latent Attribute Directions

### v1

@@ 4.4 Debiasing Multiple Attributes (intersectional debiasing)
Figure 6: Images generated from the original SD (left) and Ours (right) for gender debias with prompt ‘A photo of a work’. Gendet ratio: Male : Female = 13 : 2 $\rightarrow$ 8 : 7
Our method can be inherently extended to debiasing multiple attributes in diffusion models (DM). Specifically, in the case of multiple attribute debiasing, the adapter for each attribute should not interfere with the others. Otherwise, the most recently trained adapter could degrade the performance of previously learned adapters. For each attribute to be debiased, we denote a set of adapter parameters as $\left\{\bm{P}_{t},\bm{Q}_{t}\right\}$. We have $\bm{P}_{t}=\left[\bm{p}_{t}^{1},\bm{p}_{t}^{2},\ldots,\bm{p}_{t}^{r}\right],\bm{Q}_{t}=\left[\bm{q}_{t}^{1},\bm{q}_{t}^{2},\ldots,\bm{q}_{t}^{r}\right]$.
To avoid interference between attribute adapters, we extend Eq. 7 by introducing an orthogonal regularization loss that regularizes the vector subspace spanned by each $\bm{P}_{t}$ and $\bm{Q}_{t}$ to be orthogonal to each other:
$\mathcal{L}_{\text{orth}}=\sum_{i=1}^{t-1}\left(\bm{P}_{i}\times\bm{P}_{t}+\bm{Q}_{i}\times\bm{Q}_{t}\right).\$ (11)
$\mathcal{L}=\mathcal{L}_{\text{Guidance}}+\gamma\mathcal{L}_{\text{orth}}.$ (12)
@@ 5 Experiment
In this section, we first describe the experimental details and evaluation metrics, then present a quantitative and qualitative analysis of our method. Please refer to Appendix 9 for more results.
@@ 5.1 Experimental Details
We use Stable Diffusion v2.1 for all methods. We employ the prompt template “a photo of the face of a {occupation}, a person”. At inference time, for each bias, we generate 100 images per occupation across 100 occupations, resulting in a total of 10,000 images. We set $\eta=\alpha=1$, and train for 1000 iterations with a learning rate of 1e-5. For gender bias, we use the CelebA [12] dataset to train a binary classifier with two categories:{male,female}. For racial bias, we use the FairFace [10] dataset to train a classifier with the following four categories: WMELH={White, Middle Eastern, Latino Hispanic}, Asian={East Asian, Southeast Asian}, Black, and Indian. Please refer to supplementary for more details.
@@
In this work, we compare our method with recent state-of-the-art (SOTA) methods, including a retraining-based method, Finetuning for Fairness (F4Fair) [30], a training-free approach, H-Distribution Guidance (H Guidance) [18], a closed-form editing approach, Unified Concept Editing (UCE) [8] and an h-space manipulation approach, Interpretable Diffusion [11]. Please refer to Appendix 8 for more detailed introduction of these methods.
@@ 5.2 Evaluation Metrics
Fairness Discrepancy (FD). Following prior work[18], we adopt the Fairness Discrepancy (FD) metric. For an attribute $a$ and target distribution ${p^{a}_{\theta}}$, we use a high-accuracy classifier $\mathcal{C}_{a}$ to compute the fairness performance: $||{p^{a}_{\theta}}-\mathbb{E}_{\mathbf{x}\sim p_{\theta}(\mathbf{x})}(\mathbf{y})||_{2}$ where $\mathbf{y}$ is the softmax output of $\mathcal{C}_{a}(\mathbf{x})$. The target distribution ${p^{a}_{\theta}}$ can be any user-defined vector, typically uniform. A lower FD score indicates a closer match to the target distribution.
@@
BRISQUE. The generated image quality is also important as the debiasing process should not influence the image generation ability. We use the BRISQUE metric for evaluate the quality of the generated images as in the prior work [18].
Figure 7: Images generated from original SD (left) and Ours (right) for race debias with prompt ‘A photo of a sportsman’. Racial group distribution: WMELH : Asian : Black:Indian = 7:2:5:1 $\rightarrow$ 4:4:4:3
@@ 5.3 Results
Table 1 demonstrates that our method outperforms others in mitigating gender bias over two predefined distributions: ${f^{1}_{\theta}}=(0.5,0.5)$, representing equal likelihood of male and female, and ${f^{2}_{\theta}}=(0.2,0.8)$, where male and female have a 20% and 80% probability, respectively. Original SD model exhibits high FD scores (0.424 at $f^{1}_{\theta}$ and 0.847 at $f^{2}_{\theta}$), indicating significant bias towards one gender. While previous methods like F4Fair, H Guidance, and UCE reduce bias to some extent, our method achieves the most reduction, with FD scores of 0.003 at $f^{1}_{\theta}$ and 0.005 at $f^{2}_{\theta}$. Additionally, our method preserves semantic similarity (CLIPsim of 0.38) and image quality (BRISQUE scores of 38.46 and 38.72), matching or slightly surpassing the Original SD model. The qualitative results in Figure 6 further demonstrate that our approach effectively mitigates gender bias without compromising image quality or semantic coherence.
@@
Table 2: Comparisons of our method to the SOTA methods in racial bias over two distributions: ${f^{1}_{\theta}}=(0.25,0.25,0.25,0.25)$ and ${f^{2}_{\theta}}=(0.4,0.3,0.2,0.1)$, representing probability of WMELH, Asian, Black, and Indian respectively.
@@
We further evaluate the transferability of our method by training it with Stable Diffusion v2.1 and testing it over other versions, and results are reported in Table 3. From the table we can see that the performance of in all metrics are close to the optimal results achieved when training and testing are performed using the same model version (v2.1). While there is a slight increase in FD when tested on v1.4, v1.5, and v2.0, the differences are minor, indicating that our method can effectively generalize across different model versions. This robustness highlights the method’s capability to transfer learned features across varying model conditions, underscoring its practical value for real-world applications.
@@ 6 Conclusion
In this paper, we propose DebiasDiff, a plug-and-play method that learns attribute latent directions in a self-discovering manner, thus mitigates the reliance on collecting additional reference datasets. Our method can not only jointly debias multiple attributes in DMs, but also enables the generated images to follow a prescribed attribute distribution. It is lightweight and can be integrated with other DMs without re-training. Extensive experiments on debiasing gender, racial, and their intersectional biases show that our method outperforms previous SOTA by a large margin. We believe that our work marks a critical advancement in addressing harmful societal stereotypes within diffusion models, and it contributes to the ethical real-world applications of text-to-image diffusion models.
@@ 8 Compared Methods
Finetuning for Fairness (F4Fair) [30] is a training-free approach with two main technical innovations: (1) a distributional alignment loss that aligns specific attributes of generated images to a user-defined target distribution, and (2) adjusted direct finetuning (adjusted DFT) of the diffusion model’s sampling process, which uses an adjusted gradient to directly optimize losses on generated images.
H-Distribution Guidance (H Guidance) [18] is another training-free approach. It introduces Distribution Guidance, which ensures that generated images follow a prescribed attribute distribution. This is achieved by leveraging the latent features of the denoising UNet, which contain rich demographic semantics, to guide debiased generation. They also train an Attribute Distribution Predictor (ADP), a small MLP that maps latent features to attribute distributions. ADP is trained using pseudo labels generated by existing attribute classifiers, allowing fairer generation with the proposed Distribution Guidance.

### v3

@@ 4.4 Debiasing Multiple Attributes (intersectional debiasing)
Figure 6: Images generated from the original Stable Diffusion model (left) and FairGen (right) for gender debiasing with the prompt “A photo of a worker”. The gender distribution shifts from a male-to-female ratio of 13:2 in the original model to 8:7 with FairGen.
Our method can be inherently extended to debiasing multiple attributes in diffusion models (DM). Specifically, in the case of multiple attribute debiasing, the adapter for each attribute should not interfere with the others. Otherwise, the most recently trained adapter could degrade the performance of previously learned adapters. For each attribute to be debiased, we denote a set of adapter parameters as $\left\{\boldsymbol{P}_{t},\boldsymbol{Q}_{t}\right\}$. We have $\boldsymbol{P}_{t}=\left[\boldsymbol{p}_{t}^{1},\boldsymbol{p}_{t}^{2},\ldots,\boldsymbol{p}_{t}^{r}\right],\boldsymbol{Q}_{t}=\left[\boldsymbol{q}_{t}^{1},\boldsymbol{q}_{t}^{2},\ldots,\boldsymbol{q}_{t}^{r}\right]$.
To avoid interference between attribute adapters, we extend Eq. 7 by introducing an orthogonal regularization loss that regularizes the vector subspace spanned by each $\boldsymbol{P}_{t}$ and $\boldsymbol{Q}_{t}$ to be orthogonal to each other:
$\mathcal{L}_{\text{orth}}=\sum_{i=1}^{t-1}\left(\boldsymbol{P}_{i}\times\boldsymbol{P}_{t}+\boldsymbol{Q}_{i}\times\boldsymbol{Q}_{t}\right),\\$ (11)
$\mathcal{L}=\mathcal{L}_{\text{Guidance}}+\gamma\cdot\mathcal{L}_{\text{orth}}.$ (12)
@@ 5 Experiment
In this section, we first describe the experimental details and evaluation metrics, then present a quantitative and qualitative analysis of our method. Please refer to Appendix for more results.
@@ 5.1 Experimental Details
We use Stable Diffusion v2.1 for all methods. We employ the prompt template “a photo of the face of a {occupation}, a person”. At inference time, for each bias, we generate 100 images per occupation across 100 occupations, resulting in a total of 10,000 images. We set $\eta=\alpha=0.3$, and train for 1000 iterations with a learning rate of 1e-5. For gender bias, we use the CelebA [12] dataset to train a binary classifier with two categories:{male,female}. For racial bias, we use the FairFace [10] dataset to train a classifier with the following four categories: WMELH={White, Middle Eastern, Latino Hispanic}, Asian={East Asian, Southeast Asian}, Black, and Indian. Please refer to the supplementary for more details.
@@
In this work, we compare our method with recent state-of-the-art (SOTA) methods, including two retraining-based methods, Finetuning for Fairness (F4Fair) [30] and Inclusive Text-to-Image Generation (ITI-GEN) [37], an approach that does not require retraining DM, H-Distribution Guidance (H Guidance) [18], a closed-form editing approach, Unified Concept Editing (UCE) [8] and an h-space manipulation approach, Interpretable Diffusion [11]. Please refer to Appendix for more detailed introduction of these methods.
@@ 5.2 Evaluation Metrics
Fairness Discrepancy (FD). Following prior work[18], we adopt the Fairness Discrepancy (FD) metric. For an attribute $a$ and target distribution $p^{a}_{\theta}$}, we use a high-accuracy classifier $\mathcal{C}_{a}$ to compute the fairness performance: ${p^{a}_{\theta}}-\mathbb{E}_{\mathbf{x}\sim p_{\theta}(\mathbf{x})}(\mathbf{y})||_{2}$ where $\mathbf{y}$ is the softmax output of $\mathcal{C}_{a}(\mathbf{x})$. The target distribution $p^{a}_{\theta}$} can be any user-defined vector, typically uniform. A lower FD score indicates a closer match to the target distribution.
@@
BRISQUE. The generated image quality is also important as the debiasing process should not influence the image generation ability. We use the BRISQUE metric for evaluate the quality of the generated images.
Figure 7: Images generated from the original Stable Diffusion model (left) and FairGen (right) for racial debiasing with the prompt “A photo of a sportsman”. The racial group distribution shifts from WMELH : Asian : Black : Indian = 7:2:5:1 in the original model to 4:4:4:3 with FairGen.
@@ 5.3 Results
Table 1 demonstrates that our method outperforms others in mitigating gender bias over two predefined distributions: ${f^{1}_{\theta}}=(0.5,0.5)$, representing equal likelihood of male and female, and ${f^{2}_{\theta}}=(0.2,0.8)$, where male and female have a 20% and 80% probability, respectively. Original SD model exhibits high FD scores (0.424 at $f^{1}_{\theta}$ and 0.847 at $f^{2}_{\theta}$), indicating significant bias towards one gender. While previous methods like F4Fair, ITI-GEN, H Guidance, and UCE reduce bias to some extent, our method achieves the most reduction, with FD scores of 0.003 at $f^{1}_{\theta}$ and 0.005 at $f^{2}_{\theta}$. Additionally, our method preserves semantic similarity (CLIPsim of 0.38) and image quality (BRISQUE scores of 38.46 and 38.72), matching or slightly surpassing the Original SD model. The qualitative results in Figure 6 further demonstrate that our approach effectively mitigates gender bias without compromising image quality or semantic coherence.
@@
Table 2: Comparisons of our method to the SOTA methods in racial bias over two distributions: ${f^{1}_{\theta}}=(0.25,0.25,0.25,0.25)$ and ${f^{2}_{\theta}}=(0.4,0.3,0.2,0.1)$, representing the probability of WMELH, Asian, Black, and Indian respectively.
@@
We further evaluate the transferability of our method by training it with Stable Diffusion v2.1 and testing it over other versions, and results are reported in Table 3. From the table we can see that the performance of in all metrics are close to the optimal results achieved when training and testing are performed using the same model version (v2.1). While there is a slight increase in FD when tested on v1.4, v1.5, and v2.0, the differences are minor, indicating that our method can effectively generalize across different model versions. This robustness highlights the method’s capability to transfer learned features across varying model conditions.
@@ 6 Conclusion
In this paper, we propose FairGen, a plug-and-play method that learns attribute latent directions in a self-discovering manner. Our method can not only jointly debias multiple attributes in DMs, but also enables the generated images to follow a prescribed attribute distribution. It is lightweight and can be integrated with other DMs without retraining. Extensive experiments on debiasing gender, racial, and their intersectional biases show that our method outperforms previous SOTA by a large margin. We believe that our work marks a critical advancement in addressing harmful societal stereotypes within diffusion models, and it contributes to the ethical real-world applications of text-to-image DMs.
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grant No. 62306261, and the Shun Hing Institute of Advanced Engineering (SHIAE) Fund under Grant No. 8115074.
@@ 8 Compared Methods
Finetuning for Fairness (F4Fair) [30] is a retraining-based approach with two main technical innovations: (1) a distributional alignment loss that aligns specific attributes of generated images to a user-defined target distribution, and (2) adjusted direct finetuning (adjusted DFT) of the diffusion model’s sampling process, which uses an adjusted gradient to directly optimize losses on generated images.
Inclusive Text-to-Image GENeration (ITI-GEN) [36] enhances fairness in text-to-image synthesis by incorporating reference images. Instead of relying solely on text prompts, ITI-GEN leverages visual exemplars to more effectively represent attributes that are difficult to describe in words, such as nuanced variations in skin tones. The key idea is to learn prompt embeddings that guide the generation process, ensuring balanced and inclusive outputs across different attribute categories.
H-Distribution Guidance (H Guidance) [18] does not require retraining DMs. It introduces Distribution Guidance, which ensures that generated images follow a prescribed attribute distribution. This is achieved by leveraging the latent features of the denoising UNet, which contain rich demographic semantics, to guide debiased generation. They also train an Attribute Distribution Predictor (ADP), a small MLP that maps latent features to attribute distributions. ADP is trained using pseudo labels generated by existing attribute classifiers, allowing fairer generation with the proposed Distribution Guidance.
@@ 9 More Visualization Results
10 More Results on Scalability
FairGen is designed with modularity and efficiency in mind, enabling scalable debiasing across multiple sensitive attributes. Its scalability stems from two core design choices: (1) a lightweight adapter architecture with linear complexity, and (2) an inference-time composition mechanism that avoids retraining or classifier dependency.
Linear Adapter Complexity.
For $k$ attributes with $c$ categories each, FairGen introduces $\mathcal{O}(kc)$ plug-and-play adapters. These adapters are trained independently and only the relevant ones are activated at inference time, ensuring that the computational cost remains bounded and practical, even as the number of attributes grows.
Training Efficiency.
Each adapter is trained within 0.5 hours on a single A100 GPU. Even under multi-attribute settings, the total training time remains competitive with prior methods. As shown in Table 7, FairGen achieves the lowest training time (1.0h) and identical inference speed (6.2s) to the base diffusion model, demonstrating its lightweight nature.
Table 7: Training and inference efficiency comparison.
Metric Original SD F4Fair ITI-GEN H Guidance InterDiff FairGen (Ours)
Training Time 4.3 h 2.4 h 2.8 h 3.1 h 1.0 h
Inference Time 6.2s 6.8s 6.4s 7.1s 6.9s 6.2s
Inference Efficiency.
During inference, FairGen introduces negligible overhead by only activating the adapters corresponding to the target attribute categories. As Table 7 shows, its inference time matches that of the original Stable Diffusion, while achieving superior fairness.
Scalability in Attribute Space.
To evaluate FairGen’s robustness in higher-dimensional fairness settings, we extend it to debias across four attributes: gender, race, age (“young”, “middle-aged”, “old”), and body type (“thin”, “medium”, “obese”). Table 8 shows that FairGen consistently achieves low fairness discrepancy (FD) and maintains visual quality (CLIPsim, FID, BRIS) with only minor degradation, confirming its practical scalability.
Table 8: Scalability analysis of FairGen on four attributes.
Setting FD $\downarrow$ CLIPsim $\uparrow$ FID $\downarrow$ BRIS $\uparrow$
Gender 0.041 0.37 12.34 38.52
Gender+Race 0.042 0.37 13.18 38.33
Gender+Race+Age 0.044 0.36 13.95 38.41
Gender+Race+Age+Body Type 0.045 0.36 13.78 38.27
Orthogonality Regularization.
FairGen applies orthogonality regularization to mitigate attribute interference during training. While helpful, we observe that strict orthogonality is not essential for strong debiasing. Due to the lightweight adapter structure, the additional cost introduced by this regularization remains minor. In future work, scalability can be further improved by applying orthogonality selectively, e.g., only between interfering attribute groups identified through data-driven analysis.
Overall, these results demonstrate FairGen’s ability to scale to a broad range of fairness settings while maintaining efficiency and quality, making it a practical solution for large-scale fair image generation.